# LVAgent: Long Video Understanding by Multi-Round Dynamical Collaboration of MLLM Agents

### v4

@@ 4.3 Comparison with SOTA
As shown in Table 1, we compare the methods with closed-source models [38, 44], open-source MLLMs [2, 24, 4, 47, 61, 40, 62, 22, 28, 20, 6], and agent-based systems [49, 51, 58, 33, 32]. Experimental results show that our method is the first agent-based method to achieve over 80% accuracy across these four benchmarks. Especially on LongVideoBench, LVAgent outperforms GPT-4o (SOTA) by 13.3%. On MLVU, LVAgent surpasses GPT-4o and InternVL-2.5 by 19.3% and 8.2%.
@@ 6 Acknowledgement
This work was supported by the National Key R&D Program of China (NO.2022ZD0160505). I sincerely express the gratitude for Zijun Liu on valuable discussions about multi-round dynamical collaboration of MLLM agents.

### v5

@@ 4.3 Comparison with SOTA
As shown in Table 1, we compare the methods with closed-source models [gpt4o, gemini15], open-source MLLMs [chen2024sharegpt4video, videochat, internvl2.5, qwen2vl, longva, longvu, llavavideo, aria, liu2024oryx, llavaonevision, videollama2], and agent-based systems [videoagentrepeat, videotree, vca, drvideo, videoragvideofeat]. Experimental results show that our method is the first agent-based method to achieve over 80% accuracy across these four benchmarks. Especially on LongVideoBench, LVAgent outperforms GPT-4o (SOTA) by 13.3%. On MLVU, LVAgent surpasses GPT-4o and InternVL-2.5 by 19.3% and 8.2%.
@@ 6 Acknowledgement
This work was supported by the National Key R&D Program of China (NO.2022ZD0160505). We sincerely express the gratitude for Zijun Liu on valuable discussions about multi-round dynamical collaboration of MLLM agents.